A jar contains 13 marbles, 3 red, 4 white, and 6 blue. Two marbles are drawn without replacement.What is the probability that they are the same color?
a) 4/13 b) 9/13 c) 10/13 d) None of a, b, or c.

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12569)

2 reds=3/13*2/12=6/156
2 whites=4/13*3/12=12/156
2 blues=6/13*5/12=30/156
(6+12+30)/156=48/156=4/13 ans.

Answer by stanbon(57424)

A jar contains 13 marbles, 3 red, 4 white, and 6 blue. Two marbles are drawn without replacement.What is the probability that they are the same color?
------
Prob(red or white or blue) = [3C2 + 4C2 + 6C2]/[13C2] = (3+6+15)/78 = 24/78
= 12/39 = 4/13
=============================
Cheers,
Stan H.
=============================
a) 4/13 b) 9/13 c) 10/13 d) None of a, b, or c.